# Neural Exploratory Landscape Analysis

### v2

@@ 3.2 Architecture of $\Lambda_{\theta}$
Ts-Attn. The computation graph of the Ts-Attn is illustrated on the right side of Figure 2, with a basic component being the attention block ($Attn$) of hidden dimension $h$. As illustrated in the left side of Figure 2, the $Attn$ block follows the design of the original Transformer [25], except that the layer normalization [53] is used instead of batch normalization [54]. Ts-Attn receives $E^{t}$ and then advances the information sharing at both cross-solution and cross-dimension levels. 1) Cross-solution attention: We utilize an Attn block to enable same dimensions shared by different candidate solutions within the population to exchange information. 2) Cross-dimension attention: A second Attn block is implemented to further promote the sharing of information across different dimensions within each candidate. To achieve this, the output from the cross-solution phase must be transposed into the shape $m\times d\times h$ and augmented with cosine/sine positional encodings to maintain the dimensional order within a candidate. After these two phases of information processing, the resulting output $s^{t}$ provides comprehensive optimization status features, suitable for different MetaBBO algorithms. This highly parallelizable, attention-based architecture significantly boosts the scalability of NeurELA as the number of candidates $m$ or dimensions $d$ increases. Additional technical details about the architecture of $\Lambda_{\theta}$ can be found in Appendix A.1.
@@ 3.3 Training Method
Neuroevolution of $\Lambda_{\theta}$. The meta-objective of our NeurELA, as defined in Eq. (4), is non-differentiable, since the evaluation of the relative performance $\Upsilon\left(\Lambda_{\theta}\mid\mathbb{T}_{k}\right)$ in Eq. (3) is not directly derived from the parameter ${\theta}$ of $\Lambda_{\theta}$. We in turn employ an Evolution Strategy, denoted as $ES$, to neuroevolve $\Lambda_{\theta}$ towards maximizing its expected relative performance for multiple MetaBBO tasks in $\Omega$. The pseudo code of the training is presented in Algorithm 1. The $ES$ first initializes a group of $N$ landscape analysers, then iteratively searches for ${\theta^{*}}$ through $maxGen$ generations of optimization (lines $4$ to $16$). Note that the meta-training of the $K$ MetaBBO tasks, which involves training their neural-policies over the training problem set, is nested within the optimization loop (lines $5$ to $12$) and is extremely time-consuming. We hence introduce Ray [55], an open-source framework for parallel processing in machine learning applications, to distribute the meta-training pipelines to multiple CPUs. This parallelization significantly enhances the overall efficiency of training NeurELA.
@@ 4 Experimental Results
Training Setup. The operating space $\Omega$ comprises $K=3$ MetaBBO tasks. In specific, we first select three representative MetaBBO algorithms LDE [21], RLEPSO [56]and RL-DAS [10] from existing MetaBBO literature. Then we set the associated problem set $\mathbb{D}_{k}$ for these MetaBBO algorithms as the BBOB testsuites in COCO [57], which includes a variety of optimization problems with diverse landscape properties. We have to note that the selected algorithms cover diverse MetaBBO scenarios such as auto-configuration for control parameters and auto-selection for evolutionary operators, hence ensuring the generalization of our NeurELA. We follow the original settings of these MetaBBO algorithms during their meta-training, as suugested in their papers. For the settings about the neural network, we set the hidden dimension $h=16$ for the neural network modules in the landscape analyser $\Lambda_{\theta}$. With a single-head attention in $Attn$, $\Lambda_{\theta}$ possess a total number of $3296$ learnable parameters. We adopt Fast CMA-ES [19] as $ES$ for its searching efficiency and robust optimization performance. During the training, we employ a population of $N=10$ $\Lambda_{\theta}$s within a generation and optimize the population for $maxGen=50$ generations. For each generation, we parallel the $N\times K=30$ meta-training pipelines across $30$ independent CPU cores using Ray [55]. For each low-level optimization process, we set the maximum functiona evaluations as 20000. The experiments are run on a computation node of a Slurm CPU cluster, with two Intel Sapphire Rapids 6458Q CPUs and $256$ GB memories. Due to the limitation of the space, we present more technical details such as the train-test split $\{\mathbb{D}_{\text{train}},\mathbb{D}_{\text{test}}\}$, the control parameters of Fast CMAES and the use of open-sourced software in Appendix A.2 $\sim$ A.5.
Baselines. During the following comparison experiments, we consider two baselines for comparison: 1) Original, which denotes the original hand-crafted landscape analyser functions ($\Lambda_{0}$) in the MetaBBO algorithms. 2) ELA, which denotes the traditional ELA feature groups summarized in Section 2 and defined in [15], which we implement by the open-sourced Pflacco package [39].
@@ 4.1 Zero-shot Performance (RQ1)
Recall that an MetaBBO task is composed of an MetaBBO algorithm and an associated optimization problem set (i.e., $\mathbb{T}_{k}=\left(\mathbb{A}_{k},\mathbb{D}_{k}\right)$). We hence generalize the pre-trained landscape analyser $\Lambda_{\theta^{*}}$ towards three zero-shot cases: 1) five unseen MetaBBO algorithms DEDDQN [26], RLPSO [23], GLEET [24], MelBA [58] and GLHF [31]; 2) two unseen problem set BBOB-Noisy [57] and Protein-Docking benchmark [59]; 3) both the unseen MetaBBO algorithms and problem sets. Those tested MetaBBO tasks cover a wide range of meta-learned black-box optimizers. Besides, the problem sets include challenging realistic application, e.g., Protein-Docking. We evaluate the performance of the ELA and our NeurELA by calculating the relative optimization performance (Eq. (3)) against the Original baseline in the given MetaBBO tasks. Since the performance of Original baseline would be 00 within this setting, we report the exponential of the calculated relative performances in Figure 3, where the bars denote the average relative performance, the error bars denote the $95\%$ bootstrapped confidence interval, for $10$ runs of meta-training. For BBOB-Noisy set, we set the maximum function evaluations as 20000, while for Protein-Docking set, we set it to 1000. The three zero-shot cases are split by the dashed lines.
@@
The results in Figure 3 show that: 1) Over all of the three zero-shot cases, NeurELA generally achieves superior performance against original feature extraction designs (Original) in these MetaBBO algorithms, as well as the traditional landscape analysis features (ELA). 2) The results of integrating ELA into the evaluated MetaBBO tasks are quite noisy, which might be due to the intricate correlations in-between the feature groups. Such strong correlations might challenge the information processing ability of the neural policies $\Pi_{\phi}$ in some MetaBBO algorithms. 3) We observe that our NeurELA significantly boosts the performance of Original baseline on unseen MetaBBO algorithms & problem sets case, which not only validates the effectiveness of our design, but also indicates that the specific feature extraction designs (Origin) in existing MetaBBO algorithms might overfits to the problem set it is meta-trained on. 4) Our NeurELA consistently outperforms the traditional ELA features on the BBOB-Noisy testsuites and Protein-Docking benchmark, which comprises a wide range of optimization problems with either different noise models and levels (e.g., Gaussian noise and Cauchy noise [60]) or intricate realistic problem features. It indicates that our NeurELA shows robust generalization ability to unseen MetaBBO tasks, which involve both unseen MetaBBO algorithm and unseen realistic optimization problems.
@@
Figure 4: Performance gain curves of fine-tuning NeurELA for a specific MetaBBO task. The three tasks are selected to illustrate scenarios where the pre-trained NeurELA outperforms, matches, or underperforms the original baseline.
@@ 4.2 Fine-tuning Performance (RQ2)
We select three representative MetaBBO task from the zero-shot performance analysis (Figure 3) to validate fine-tuning adaption efficiency of our NeurELA. The selected MetaBBO tasks includes: 1) LDE on BBOB problem set, where NeurELA surpasses the Original baseline in LDE. 2) RLPSO on Protein-Docking problem set, where NeurELA equally performs against the Original baseline in RLPSO. 3) GLEET on BBOB problem set, where NeurELA does not achieves the expected zero-shot performance. Concretely, we integrate the pre-trained $\Lambda_{\theta^{*}}$ into the corresponding MetaBBO algorithm, then fine-tune it in conjunction with the meta-training over the given problem set. We depict the performance gain curves during the meta-training in Figure 4. We additionally plot the performance gain curves of the meta-training during the zero-shot generalization for comparison. The results showcase the flexible adaption potential of our NeurELA: 1) fine-tuning NeurELA on specific MetaBBO tasks where our NeurELA outperforms, equally performs or underperforms against the Original baseline would consistently introduce significant performance boosts. 2) the fine-tuning strategy shows remarkable efficiency in adapting NeurELA for new MetaBBO algorithm dynamics and optimization problem structures, which only requires less than $20\%$ meta-training epochs to achieve the similar zero-shot performance.
@@ 4.3 What has NeurELA learned? (RQ3)
In this section, we presents the advantage of the landscape features learned by NeurELA against traditional ELA features under the MetaBBO settings. Concretely, we take LDE [21] as a showcase. In LDE, the meta-level policy (a LSTM-based RL agent) outputs the mutation strength factor $F^{t}\in(0,1)$ for the low-level DE per optimization step $t$. $F^{t}$ indicates the exploration-exploitation tradeoff during the optimization (larger value indicates exploration, vice versa). We first record the per step populations during the optimization process of LDE ($10$ runs on 10-D Rastrigin in CoCo-BBOB testsuites). We then simply classify the recorded populations as exploration population ($F^{t}>0.5$) and exploitation population ($F^{t}\leq 0.5$). We now illustrate the low-rank visualization of the ELA feature and our NeurELA feature for the recorded populations in Figure 5. The data points denoted the 2-dimensional features obtained by first feeding the populations into ELA/NeurELA and then using PCA [61] transformation on the output landscape features. Red points and blue points represent the features for the exploration and exploitation population respectively. The results validate that NeurELA is more sensitive than traditional ELA features on the dynamics of low-level optimization. Since traditional ELA is mainly used for describing an optimization problem, it might falls short in providing informative low-level optimization status for MetaBBO methods. This explains the consistent superior performance of our NeurELA (see Figure 3). Compared with the traditional ELA features, our NeurELA can be meta-trained on MetaBBO tasks to provide sensitive and accurate low-level optimization status for the meta-level policy. As shown in Figure 5, the decision bound between the exploration population and the exploitation population is clear, which is helpful for the meta-level policy to balance the exploration-exploitation tradeoff within the low-level optimization.
@@ 4.4 In-depth Study (RQ4 $\sim$ RQ6)
Computational Efficiency Comparison (RQ3). We compare the computational efficiency of our NeurELA with Original and ELA as either the amount of candidate samples $m$ or the dimension of the optimization problem $d$ scales. We report the average wall-time required by each baseline and our NeurELA to produce the landscape features (for $1000$ runs). The results show that: 1) Our NeurELA achieves comparative computational efficiency with the specific hand-crafted designs in existing MetaBBO algorithms, while introducing significant performance boosts; 2) When the sample size $m$ increases, our NeurELA becomes more time-efficient than the original feature extraction designs in existing MetaBBO algorithms. This is primarily owning to the two-stage attention architecture in our NeurELA, which can be efficiently computed on GPUs in parallel. 3) The traditional ELA features encounters the curse of dimensionality, which makes itself impractical for MetaBBO algorithms to solve high-dimensional optimization problems.
Evolution Strategy Selection (RQ4). Within this paper, we employ a neuroevolution process to optimize the parameters of the landscape analyzer $\Lambda_{\theta}$, which presents a large-scale optimization challenge as $\Lambda_{\theta}$ comprises $3296$ learnable parameters. To identify an effective $ES$, we implement several candidate evolution strategy variants: Fast CMAES [19], SEP-CMAES [62], R1ES, RMES [63], and the original CMAES [18], and report the overall zero-shot performance (unseen algorithms & unseen problem sets) achieved by training our NeurELA with these candidate strategies for $10$ runs in Figure 7. The results reveal that Fast CMAES dominates the other candidates, and is hence selected as the neural network optimizer $ES$ for our NeurELA.
@@
Model Complexity (RQ5). We discuss the relationship between the model complexity and the zero-shot performance (unseen MetaBBO algorithm & problem sets) of our NeurELA. Concretely, We pre-train NeurELA under $6$ different model complexities, with various hidden dimensions, i.e., $h=(16,64)$, and the number of the Ts-Attn module, i.e., $l=(1,3,5)$. We additionally pre-train three MLP baselines, which substitute the Ts-Attn module in NeurELA with a linear feed-forward layer, which holds a shape of $h\times h,h=(16,64,128)$. We report both the zero-shot performance ($y$-axis) and the computational efficiency ($x$-axis, presented as the consumed wall time for computing the landscape features) in Figure 7, where the dashed lines denotes the performance and wall-time of the Original baseline. #$\#para$ denotes the number of the learnable parameters. The results show that: 1) a significant performance gap is observed between the MLP baselines and our Ts-Attn module ($h=16,l=1$). It validates the effectiveness of our Ts-Attn design, which enhance the feature extraction of our NeurELA by encouraging the information sharing at both the cross-solution and cross-dimension levels; 2) As the model complexity increases, the performance of the Ts-Attn module drops rapidly. It reveals that the increased number of learnable parameters challenges the optimization ability of the backend $ES$. Given the limited computational resources, it is difficult to identify the optimal parameters ${\theta^{*}}$.

### v3

@@ 3.2 Architecture of $\Lambda_{\theta}$
Ts-Attn. The computation graph of the Ts-Attn is illustrated on the right side of Figure 2, with a basic component being the attention block ($Attn$) of hidden dimension $h$. As illustrated in the left side of Figure 2, the $Attn$ block follows the design of the original Transformer Vaswani et al. (2017), except that the layer normalization Ba et al. (2016) is used instead of batch normalization Ioffe & Szegedy (2015). Ts-Attn receives $E^{t}$ and then advances the information sharing at both cross-solution and cross-dimension levels. 1) Cross-solution attention: We utilize an Attn block to enable same dimensions shared by different candidate solutions within the population to exchange information. 2) Cross-dimension attention: A second Attn block is implemented to further promote the sharing of information across different dimensions within each candidate. To achieve this, the output from the cross-solution phase must be transposed into the shape $m\times d\times h$ and augmented with cosine/sine positional encodings to maintain the dimensional order within a candidate. After these two phases of information processing, the resulting output $s^{t}$ provides comprehensive optimization status features, suitable for different MetaBBO algorithms. This highly parallelizable, attention-based architecture significantly boosts the scalability of NeurELA as the number of candidates $m$ or dimensions $d$ increases. Additional technical details about the architecture of $\Lambda_{\theta}$ can be found in Appendix A.1.
@@ 3.3 Training Method
Neuroevolution of $\Lambda_{\theta}$. The meta-objective of our NeurELA, as defined in Eq. (4), is non-differentiable, since the evaluation of the relative performance $\Upsilon\left(\Lambda_{\theta}\mid\mathbb{T}_{k}\right)$ in Eq. (3) is not directly derived from the parameter ${\theta}$ of $\Lambda_{\theta}$. We in turn employ an Evolution Strategy, denoted as $ES$, to neuroevolve $\Lambda_{\theta}$ towards maximizing its expected relative performance for multiple MetaBBO tasks in $\Omega$. The pseudo code of the training is presented in Algorithm 1. The $ES$ first initializes a group of $N$ landscape analysers, then iteratively searches for ${\theta^{*}}$ through $maxGen$ generations of optimization (lines $4$ to $16$). Note that the meta-training of the $K$ MetaBBO tasks, which involves training their neural-policies over the training problem set, is nested within the optimization loop (lines $5$ to $12$) and is extremely time-consuming. We hence employ Ray (Moritz et al., 2018), an open-source framework for parallel processing in machine learning applications, to distribute the meta-training pipelines to multiple CPUs. This parallelization significantly enhances the overall efficiency of training NeurELA.
@@ 4 Experimental Results
Training Setup. The operating space $\Omega$ comprises $K=3$ MetaBBO tasks. In specific, we first select three representative MetaBBO algorithms LDE Sun et al. (2021), RLEPSO Yin et al. (2021)and RL-DAS Guo et al. (2024a) from existing MetaBBO literature. Then we set the associated problem set $\mathbb{D}_{k}$ for these MetaBBO algorithms as the BBOB testsuites in COCO Hansen et al. (2021), which includes a variety of optimization problems with diverse landscape properties. We have to note that the selected algorithms cover diverse MetaBBO scenarios such as auto-configuration for control parameters and auto-selection for evolutionary operators, hence ensuring the generalization of our NeurELA. We follow the original settings of these MetaBBO algorithms during their meta-training, as suugested in their papers. For the settings about the neural network, we set the hidden dimension $h=16$ for the neural network modules in the landscape analyser $\Lambda_{\theta}$. With a single-head attention in $Attn$, $\Lambda_{\theta}$ possess a total number of $3296$ learnable parameters. We adopt Fast CMA-ES Li et al. (2018) as $ES$ for its searching efficiency and robust optimization performance. During the training, we employ a population of $N=10$ $\Lambda_{\theta}$s within a generation and optimize the population for $maxGen=50$ generations. For each generation, we parallel the $N\times K=30$ meta-training pipelines across $30$ independent CPU cores using Ray Moritz et al. (2018). For each low-level optimization process, we set the maximum function evaluations as 20000. The experiments are run on a computation node of a Slurm CPU cluster, with two Intel Sapphire Rapids 6458Q CPUs and $256$ GB memories. Due to the limitation of the space, we present more technical details such as the train-test split $\{\mathbb{D}_{\text{train}},\mathbb{D}_{\text{test}}\}$, the control parameters of Fast CMAES and the use of open-sourced software in Appendix A.2 $\sim$ A.5.
Baselines. We consider three baselines for comparison: 1) Original, which denotes the original hand-crafted landscape analyser functions ($\Lambda_{0}$) in the MetaBBO algorithms. 2) ELA, which denotes the traditional ELA feature groups summarized in Section 2 and defined in (Mersmann et al., 2011), which we implement by the open-sourced Pflacco package (Kerschke & Trautmann, 2019b). 3) DeepELA (Seiler et al., 2024), which is a recent work which trains an attention-based neural network as an alternative of traditional ELA features through self-supervised contrastive learning.
@@ 4.1 Zero-shot Performance (RQ1)
Recall that a MetaBBO task is composed of an MetaBBO algorithm and an associated optimization problem set (i.e., $\mathbb{T}_{k}=\left(\mathbb{A}_{k},\mathbb{D}_{k}\right)$). We hence generalize the pre-trained landscape analyser $\Lambda_{\theta^{*}}$ towards three zero-shot cases: 1) five unseen MetaBBO algorithms DEDDQN Sharma et al. (2019), RLPSO Wu & Wang (2022), GLEET Ma et al. (2024), MelBA (Chaybouti et al., 2022) and GLHF (Li et al., 2024); 2) two unseen problem set BBOB-Noisy Hansen et al. (2021) and Protein-Docking benchmark Hwang et al. (2010); 3) both the unseen MetaBBO algorithms and problem sets. Those tested MetaBBO tasks cover a wide range of meta-learned black-box optimizers. Besides, the problem sets include challenging realistic application, e.g., Protein-Docking. We evaluate the performance of the ELA, DeepELA and our NeurELA by calculating the relative optimization performance (Eq. (3)) against the Original baseline in the given MetaBBO tasks. Since the performance of Original baseline would be 00 within this setting, we report the exponential of the calculated relative performances in Figure 3, where the bars denote the average relative performance, the error bars denote the $95\%$ bootstrapped confidence interval, for $10$ runs of meta-training. For BBOB-Noisy set, we set the maximum function evaluations as 20000, while for Protein-Docking set, we set it to 1000. The three zero-shot cases are split by the dashed lines. The results in Figure 3 show that: 1) Over all of the three zero-shot cases, NeurELA generally achieves superior performance against original feature extraction designs (Original) in these MetaBBO algorithms, as well as the traditional landscape analysis features (ELA) and DeepELA. 2) The results of integrating ELA into the evaluated MetaBBO tasks are quite noisy, which might be due to the intricate correlations in-between the feature groups. Such strong correlations might challenge the information processing ability of the neural policies $\Pi_{\phi}$ in some MetaBBO algorithms. 3) We observe that our NeurELA significantly boosts the performance of Original baseline on unseen MetaBBO algorithms & problem sets case, which not only validates the effectiveness of our design, but also indicates that the specific feature extraction designs (Origin) in existing MetaBBO algorithms might overfits to the problem set it is meta-trained on. 4) Compared with traditional ELA, DeepELA show similar performance, which is in accordance with its the training objective to serve as automated alternative of traditional ELA features. 5) Our NeurELA consistently outperforms the traditional ELA features and DeepELA on the BBOB-Noisy testsuites and Protein-Docking benchmark, which comprises a wide range of optimization problems with either different noise models and levels (e.g., Gaussian noise and Cauchy noise Hansen et al. (2009)) or intricate realistic problem features. It indicates that NeurELA shows robust generalization ability to unseen MetaBBO tasks, which involve unseen MetaBBO algorithm & realistic problems.
@@
Figure 4: Performance gain curves of fine-tuning NeurELA for a specific MetaBBO task.
@@ 4.2 Fine-tuning Performance (RQ2)
We select three representative MetaBBO task from the zero-shot performance analysis (Figure 3) to validate fine-tuning adaption efficiency of our NeurELA. The selected MetaBBO tasks includes: 1) LDE on BBOB problem set, where NeurELA surpasses the Original baseline in LDE. 2) RLPSO on Protein-Docking problem set, where NeurELA equally performs against the Original baseline in RLPSO. 3) GLEET on BBOB problem set, where NeurELA does not achieves the expected zero-shot performance. Concretely, we integrate the pre-trained $\Lambda_{\theta^{*}}$ into the corresponding MetaBBO algorithm, then fine-tune it in conjunction with the meta-training over the given problem set. We depict the performance gain curves during the meta-training in Figure 4. We additionally plot the performance gain curves of the meta-training during the zero-shot generalization for comparison. The results showcase the flexible adaption potential of our NeurELA: 1) fine-tuning NeurELA on specific MetaBBO tasks where our NeurELA outperforms, equally performs or underperforms Original baseline would consistently introduce significant performance boosts. 2) the fine-tuning strategy is efficient in adapting NeurELA for new MetaBBO dynamics and problem structures, which only requires less than $20\%$ meta-training epochs to achieve the similar zero-shot performance.
@@ 4.3 What has NeurELA learned? (RQ3)
In this section, we presents the advantage of the landscape features learned by NeurELA against traditional ELA features under the MetaBBO settings. Concretely, we take LDE (Sun et al., 2021) as a showcase. In LDE, the meta-level policy (a LSTM-based RL agent) outputs the mutation strength factor $F^{t}\in(0,1)$ for the low-level DE per optimization step $t$. $F^{t}$ indicates the exploration-exploitation tradeoff during the optimization (larger value indicates exploration, vice versa). We first record the per step populations during the optimization process of LDE ($10$ runs on 10-D Rastrigin in CoCo-BBOB testsuites). We then simply classify the recorded populations as exploration population ($F^{t}>0.5$) and exploitation population ($F^{t}\leq 0.5$). We now illustrate the low-rank visualization of the ELA feature and our NeurELA feature for the recorded populations in Figure 5. The data points denoted the 2-dimensional features obtained by first feeding the populations into ELA/NeurELA and then using PCA (Shlens, 2014) transformation on the output landscape features. Red points and blue points represent the features for the exploration and exploitation population respectively. The results validate that NeurELA is more sensitive than traditional ELA features on low-level optimization dynamics. Traditional ELA is mainly used for describing a problem, it might falls short in providing informative low-level optimization status for MetaBBO methods. This explains the consistent superior performance of our NeurELA (see Figure 3). Notably, our NeurELA can be meta-trained on MetaBBO tasks to provide sensitive and accurate low-level optimization status for the meta-level policy. We also provide a further interpretation analysis which reveals the correlation of features learned by NeurELA and traditional ELA features, refer to Appendix B.3 for detailed discussion.
@@ 4.4 In-depth Study (RQ4 $\sim$ RQ6)
Computational Efficiency Comparison (RQ4). We compare the computational efficiency of our NeurELA with Original and ELA as either the amount of candidate samples $m$ or the dimension of the optimization problem $d$ scales. We report the average wall-time required by each baseline and our NeurELA to produce the landscape features (for $1000$ runs). The results show that: 1) Our NeurELA achieves comparative computational efficiency with the specific hand-crafted designs in existing MetaBBO algorithms, while introducing significant performance boosts; 2) When the sample size $m$ increases, our NeurELA becomes more time-efficient than the original feature extraction designs in existing MetaBBO algorithms. This is primarily owning to the two-stage attention architecture in our NeurELA, which can be efficiently computed on CPUs in parallel. 3) The traditional ELA features encounters the curse of dimensionality, which makes itself impractical for MetaBBO algorithms to solve high-dimensional optimization problems.
@@
Evolution Strategy Selection (RQ5). Within this paper, we employ a neuroevolution process to optimize the parameters of the landscape analyzer $\Lambda_{\theta}$, which presents a large-scale optimization challenge as $\Lambda_{\theta}$ comprises $3296$ learnable parameters. To identify an effective $ES$, we implement several candidate evolution strategy variants: Fast CMAES Li et al. (2018), SEP-CMAES Ros & Hansen (2008), R1ES, RMES Li & Zhang (2017), and the original CMAES Hansen & Ostermeier (2001b), and report the overall zero-shot performance (unseen algorithms & unseen problem sets) achieved by training our NeurELA with these candidate strategies for $10$ runs in Figure 7. The results reveal that Fast CMAES dominates the other candidates, and is hence selected as the neural network optimizer $ES$ for our NeurELA. The convergence of these ES baselines are shown in Appendix B.1.
Model Complexity (RQ6). We discuss the relationship between the model complexity and the zero-shot performance (unseen MetaBBO algorithm & problem sets) of our NeurELA. Concretely, We pre-train NeurELA under $6$ different model complexities, with various hidden dimensions, i.e., $h=(16,64)$, and the number of the Ts-Attn module, i.e., $l=(1,3,5)$. We additionally pre-train three MLP baselines, which substitute the Ts-Attn module in NeurELA with a linear feed-forward layer, which holds a shape of $h\times h,h=(16,64,128)$. We report both the zero-shot performance ($y$-axis) and the computational efficiency ($x$-axis, presented as the consumed wall time for computing the landscape features) in Figure 7, where the dashed lines denotes the performance and wall-time of the Original baseline. #$\#para$ denotes the number of the learnable parameters. The results show that: 1) a significant performance gap is observed between the MLP baselines and our Ts-Attn module ($h=16,l=1$). It validates the effectiveness of our Ts-Attn design, which enhance the feature extraction of our NeurELA by encouraging the information sharing at both the cross-solution and cross-dimension levels; 2) As the model complexity increases, the performance of the Ts-Attn module drops rapidly. It reveals that the increased number of learnable parameters challenges the optimization ability of the backend $ES$. Given the limited computational resources, it is difficult to identify the optimal parameters ${\theta^{*}}$.